# Costless correction of chain based nested sampling parameter estimation in gravitational wave data and beyond

### v1

@@ Abstract
Nested sampling parameter estimation differs from evidence estimation, in that it incurs an additional source of error. This error affects estimates of parameter means and credible intervals in gravitational wave analyses and beyond, and yet, it is typically not accounted for in standard error estimation methods. In this paper, we present two novel methods to quantify this error more accurately for any chain based nested sampler, using the additional likelihood calls made at runtime in producing independent samples. Using injected signals of black hole binary coalescences as an example, we first show concretely that the usual error estimation method is insufficient to capture the true error bar on parameter estimates. We then demonstrate how the extra points in the chains of chain based samplers may be carefully utilised to estimate this error correctly, and provide a way to check the accuracy of the resulting error bars. Finally, we discuss how this error affects $p$-$p$ plots and coverage assessments.
@@
nested sampling, parameter estimation, gravitational waves
pubyear: 2024pagerange: Costless correction of chain based nested sampling parameter estimation in gravitational wave data and beyondCostless correction of chain based nested sampling parameter estimation in gravitational wave data and beyond
@@ 1 Introduction
Nested sampling distinguishes itself from Markov Chain Monte Carlo (MCMC) algorithms (MacKay, 2003; Foreman-Mackey et al., 2013; Ashton & Talbot, 2021) in its ability to easily calculate evidences, and the errors associated with this evidence estimation are well understood (Skilling, 2006). Whilst it is also popularly used for parameter inference in gravitational wave data analysis and beyond, nested sampling parameter estimation contains an additional source of error which affects its performance (Chopin & Robert, 2010). Quantifying this error accurately is key, as it affects our estimates of parameter means and credible intervals on parameters. This error is essentially a result of the stochastic nature of the nested sampling algorithm, and so we are able to accurately quantify it by performing many nested sampling runs on the same data. However, this is computationally expensive and, for many gravitational wave problems, simply not viable. There is, as yet, little literature exploring accurately estimating this error from a single run. In response to this, (Higson et al., 2018a) proposed an innovative method of deconstructing nested sampling runs into single live point runs and using bootstrapping to determine a new error bar on parameter inferences.
Here, we present two novel approaches which utilise the extra likelihood evaluations already performed at runtime for any chain based nested sampler. Deemed to be too correlated to use in evidence estimation, the potential performance gains to parameter inference from this wealth of additional likelihood calls has been largely unexplored. As an example of this, we show that, by carefully harnessing these additional evaluations, we can correctly quantify parameter estimation errors and verify that these match the errors obtained from repeated nested sampling runs.
In the following section, we lay out necessary background on the nested sampling algorithm and identify the dominant sources of error for both evidence and parameter estimation. In Section 3, we introduce the two methods for estimating parameter inference error bars, in the context of a simulated black hole binary (BBH) signal, and demonstrate that both methods are able to accurately capture the error bar on the parameter means. We show that these methods also produce comparable results to the method presented in Higson et al. (2018a) in Section 4. How this error extends to estimates of credible intervals and coverage plots is discussed in Section 5 and, finally, our conclusions are presented in Section 6.
@@ 2.1 Anatomy of a nested sampling run
[…]
@@
Figure 1: Schematic of a typical nested sampling run with a chain based sampler. At the end of the run, we are left with a set of dead points (black), which define a series of nested iso-likelihood contours. To generate a new live point from a given point, chain based samplers use a Markov-Chain based procedure to continually generate points within its likelihood contour, until the new point is deemed uncorrelated enough with the original point from which it was seeded. Thus, at the end of the run we are also left a set of ‘phantom points’ (red), the exact number of which depends on the chain lengths. An example chain is shown in black between dead points $2$ and $3$. In parameter estimation, we typically only use the dead points and, therefore, must use the parameter values of each dead point as a proxy for the average parameter value along the entire contour. For an example two-parameter case, where the parameter being estimated is the sum $f(\theta)=\theta_{1}+\theta_{2}$, the contours of constant of constant $f(\theta)$ are shown (dashed). In this case, the parameter value of the dead points is not necessarily representative of the average parameter value over the contours, and this will be the dominant source of error in our $f(\theta)$ estimate. However, the phantom points can provide a better understanding of the variation of this parameter along the contours, enabling a more accurate quantification of this error.
@@ 2.2 Evidence estimation
In the subsequent sections of the background, we will describe the different sources of error present in evidence estimation and parameter estimation from a nested sampling run. Much of this argument follows (Higson et al., 2018a).
@@
[…]
In a typical nested sampling run, we only have access to information about a single point on each likelihood contour, not the entire contour itself. So, in practice, this integral is approximated by a sum over the dead points:
[…]
Along a given contour, by construction, all points would have the same likelihood value, so using the likelihood value of a single dead point as a proxy for $\mathcal{L}(X)$ is an exact substitution that introduces no extra error. However, the exact values of the fractional prior volume ‘shells’, $\Delta X_{i}$, are unknown, since the exact set of shrinkage ratios, $\textbf{t}_{i}$, are unknown, and this is the dominant source of error in evidence estimation. Typically, the resulting error bar on the evidence is quantified by simulating sets of $\Delta X_{i}$ to use in the evidence calculation and quoting the error from this set of estimates.
@@ 2.3 Parameter estimation
[…]
$\tilde{f}(X)$ represents the prior expectation of $f(\theta)$ given than $\mathcal{L}(\theta)=\mathcal{L}(X)$:
$\tilde{f}(X)=E^{\pi}[f(\theta)|\mathcal{L}(\theta)=\mathcal{L}(X)].$ (7)
@@
[…]
@@
Again, the exact prior volumes are unknown, but there is an additional source of error here which was not present before: given that we only have information about a single $f(\theta)$ value on each contour, we are required to use $f(\theta_{i})$ as a proxy for $\tilde{f}(X_{i})$. Unlike with the likelihood, it is no longer true that this is an exact substitution, and in parameter estimation this becomes the dominant source of error. The stochasticity of nested sampling means that over a single run we will not necessarily be able to get representative values of $f(\theta)$ along every contour, and thus won’t be able to capture our true error bar over multiple runs, even when simulating sets of prior volumes. This is demonstrated in Figure 1, where we show the example of trying to estimate the sum, $f(\theta)=\theta_{1}+\theta_{2}$, of parameters in a simple two-parameter case; here, the values of $f(\theta)$ at each of the dead points on the contours do not reflect the average value over the contours. The key to capturing this error is better understanding the variation of $f(\theta)$ along the contours.
3 Correcting parameter estimation errors with phantom points
@@
In order to study the nested sampling parameter estimation error, we consider the example of a simulated black hole binary (BBH) signal with parameters similar to those of GW150914 and with similar noise. The exact injected parameter values are shown in Figure 2. Since studying the error bars carefully requires hundreds of nested sampling runs to be performed, we work within a simplified framework in which only four of the parameters are sampled: the chirp mass ($\mathcal{M}$), mass ratio ($q$), luminosity distance ($d_{L}$), and zenith angle between the total angular momentum and line of sight ($\theta_{jn}$). For all of the below results, we perform nested sampling runs with $500$ live points per run and the default PolyChord chain length of $5*\textrm{ndims}=20$. The runs are performed using bilby (Ashton et al., 2019), with a modified version of the in-built PolyChord sampler option. All waveforms are both injected and sampled with the IMRPhenomPv2 waveform model (Hannam et al., 2014) and we use two interferometers, LIGO-Hanford (H1) and LIGO-Livingston (L1).
@@
Figure 2: Posteriors recovered from the injected signal using PolyChord, with the injected parameter values indicated. Since parameter estimation studies require large numbers of runs, we only sample these $4$ parameters, with the other parameters simply set to their injected values.
3.1 Evidence error estimation method is insufficient for parameter estimation
@@
Figure 3: Typically, a distribution of evidence estimates is computed from a single run using the ‘simulated weights’ method (top left). The mean and standard deviation of this distribution are then quoted as the evidence value and its corresponding error (black). If the error bars from single runs accurately quantify the error on the evidence due to unknown prior volumes, the overall mean evidence, computed over many runs, should lie in the $1^{\textrm{st}}$ percentile of estimates $1\%$ of the time, in the $50^{\textrm{th}}$ percentile $50\%$ of the time, and so on. Hence, plotting a histogram of the percentiles from each run in which the overall mean (blue dashed line) lies should give a uniform distribution from 0 to 100 (bottom left). We see here that this is indeed the case for the evidences, with the Kolmogorov-Smirnov $p$-value being 0.25, demonstrating that resampling the prior volumes is sufficient for estimating the error bar on the evidence for this example. For the chirp mass, even by eye the estimates per run are not wide enough to capture the variation in estimates between runs (top and middle right). This indicates that the variation in chirp mass along a given contour is not being properly accounted for. The percentiles plot (bottom right) confirms that the error estimate on a single nested sampling run is too optimistic. More often than would be the case for correctly distributed errors, the overall mean estimate for the chirp mass over 100 runs (blue dashed line) lies deep in the tails of the estimated chirp mass from a single run. The simulated weights method is not sufficient for calculating the parameter estimation error.
@@
The stochasticity of the fractional prior volumes in nested sampling means that the estimated mean evidence will vary from run to run. In order to accurately quantify the error on a single run, therefore, the error bar must reflect this run-to-run variation. For evidence estimation, this is typically achieved by sampling many sets of the shrinkage ratios, t, from a known distribution and using these to calculate the evidences, quoting the error bar from the spread of these estimates. Higson et al. (2018a) term this the ‘simulated weights method’ and we shall refer to it henceforth as such too.
@@
In the context of our simulated BBH signal, we can carefully verify that this method does indeed quantify the evidence error correctly. To do this, we first perform $100$ nested sampling runs on our simulated signal. For each run, we apply the simulated weights method using anesthetic (Handley, 2019) to calculate a set of evidence estimates. The resulting distributions of evidences are plotted in Figure 3. The dashed line represents the overall mean evidence calculated from all $100$ runs.
@@
The errors should be normally distributed, meaning that about $68\%$ of the time our ‘true evidence’ (estimated from the mean of $100$ runs) should lie within the error bars, and $95\%$ of the time it should lie within two times the extent of the error bars. This condition can also be viewed in terms of the distribution of the percentiles of the ‘true evidence’. If the errors are indeed correctly distributed, a histogram of these percentiles will yield a uniform distribution from 00 to $100$. For our simulated BBH example, this histogram is plotted in Figure 3, along with the Kolmogorov-Smirnov (K-S) test $p$-value for a uniform distribution. We can conclude from this that the simulated weights method leads to correct evidence error estimates from a single run.
@@
Skilling (2006) suggests using the same method to estimate parameter estimation errors. As an example, we shall attempt to estimate the chirp mass of our signal. As before, for each of the $100$ runs, we can simulate sets of shrinkage ratios to obtains sets of $\Delta X$ and substitute these into equation 8 to calculate a set of chirp mass estimates. These are plotted in Figure 3, along with the overall mean chirp mass estimated from all $100$ runs.
@@
This time, even by eye it seems that the resulting distributions of chirp masses per run are not quite wide enough to capture the run-to-run variation. Performing a similar analysis of the percentiles as for the evidences yields the histogram in Figure 3. It is unmistakable from both the plot and the K-S test $p$-value that the simulated weights method alone does not lead to correct error estimates on the chirp mass. Far more often than ought to be the case, the mean chirp mass over many runs lies deep in the tails of the estimates from a single run. This is an empirical verification, in the context of gravitational waves, of results already established theoretically by Chopin & Robert (2010) and experimentally by Higson et al. (2018a).
@@ 3.2 Phantom points
The key piece lacking from the simulated weights method is that it uses the single value of $f(\theta)$ available for each iso-likelihood contour as a proxy for the average $f(\theta)$ over the contour, since there is no extra information in the live points to do otherwise. However, phantom points, drawn from the likelihood-constrained prior and each with a corresponding likelihood evaluation, have the potential to provide this additional information about the variation of $f(\theta)$ along contours. Below, we present two novel error estimation methods which incorporate this, and demonstrate that they result in the correct error distributions.
3.3 Likelihood binning method

### v2

@@ Abstract
Nested sampling parameter estimation differs from evidence estimation, in that it incurs an additional source of uncertainty. This uncertainty affects estimates of parameter means and credible intervals in gravitational wave analyses and beyond, and yet, it is typically not accounted for in standard uncertainty estimation methods. In this paper, we present two novel methods to quantify this uncertainty more accurately for any chain based nested sampler, using the additional likelihood calls made at runtime in producing independent samples. Using injected signals of black hole binary coalescences as an example, we first show concretely that the usual uncertainty estimation method is insufficient to capture the true error bar on parameter estimates. We then demonstrate how the extra points in the chains of chain based samplers may be carefully utilised to estimate this uncertainty correctly, and provide a way to check the accuracy of the resulting error bars. Finally, we discuss how this uncertainty affects $p$-$p$ plots and coverage assessments.
@@
gravitational waves, methods: data analysis, methods: statistical
pubyear: 2024pagerange: Costless correction of chain based nested sampling parameter estimation in gravitational wave data and beyondB
@@ 1 Introduction
Nested sampling distinguishes itself from Markov Chain Monte Carlo (MCMC) algorithms (MacKay, 2003; Foreman-Mackey et al., 2013; Ashton & Talbot, 2021) in its ability to easily calculate evidences, and the uncertainties associated with this evidence estimation are well understood (Skilling, 2006). Whilst it is also popularly used for parameter inference in gravitational wave data analysis and beyond, nested sampling parameter estimation contains an additional source of uncertainty which affects its performance (Chopin & Robert, 2010). Quantifying this uncertainty accurately is key, as it affects our estimates of parameter means and credible intervals on parameters. This uncertainty is essentially a result of the stochastic nature of the nested sampling algorithm, and so we are able to accurately quantify it by performing many nested sampling runs on the same data. However, this is computationally expensive and, for many gravitational wave problems, simply not viable. There is, as yet, little literature exploring accurately estimating this uncertainty from a single run. In response to this, Higson et al. (2018a) proposed an innovative method of deconstructing nested sampling runs into single live point runs and using bootstrapping to determine a new error bar on parameter inferences.
Here, we present two novel approaches which utilise the extra likelihood evaluations already performed at runtime for any chain based nested sampler. Deemed to be too correlated to use in evidence estimation, the potential performance gains to parameter inference from this wealth of additional likelihood calls has been largely unexplored. As an example of this, we show that, by carefully harnessing these additional evaluations, we can correctly quantify parameter estimation uncertainties and verify that these match the uncertainties obtained from repeated nested sampling runs.
In the following section, we lay out necessary background on the nested sampling algorithm and identify the dominant sources of uncertainty for both evidence and parameter estimation. In Section 3, we first verify empirically, in the context of a simulated black hole binary (BBH) signal, that the standard methods of uncertainty estimation in nested sampling do not accurately capture the true error bars on parameter estimates. We then introduce the two new methods. Section 4 demonstrates the application of these methods to infer the correct uncertainties on parameter means and credible intervals, and show that they produce comparable results to the method presented in Higson et al. (2018a). Finally, our conclusions are presented in Section 5.
@@ 2.1 Anatomy of a nested sampling run
$X(\mathcal{L}_{i})\equiv\int_{\mathcal{L}(\theta)>\mathcal{L}_{i}}\pi(\theta)d\theta.$ (1)
@@
Figure 1: Schematic of a typical nested sampling run with a chain based sampler. At the end of the run, we are left with a set of dead points (black), which define a series of nested iso-likelihood contours. To generate a new live point from a given point, chain based samplers use a Markov-Chain based procedure to continually generate points within its likelihood contour, until the new point is deemed uncorrelated enough with the original point from which it was seeded. Thus, at the end of the run we are also left a set of ‘phantom points’ (red), the exact number of which depends on the chain lengths. An example chain is shown in black between dead points $2$ and 33. In parameter estimation, we typically only use the dead points and, therefore, must use the parameter values of each dead point as a proxy for the average parameter value along the entire contour. For an example two-parameter case, where the parameter being estimated is the sum $f(\theta)=\theta_{1}+\theta_{2}$, the contours of constant of constant $f(\theta)$ are shown (dashed). In this case, the parameter value of the dead points is not necessarily representative of the average parameter value over the contours, and this will be the dominant source of uncertainty in our $f(\theta)$ estimate. However, the phantom points can provide a better understanding of the variation of this parameter along the contours, enabling a more accurate quantification of this uncertainty.
@@ 2.2 Evidence estimation
In the subsequent sections of the background, we will describe the different sources of uncertainty present in evidence estimation and parameter estimation from a nested sampling run. Much of this argument follows Higson et al. (2018a).
@@
$\mathcal{Z}=\int_{0}^{1}\mathcal{L}(X)dX.$ (4)
In practice, this integral is approximated by a sum over the dead points:
$\mathcal{Z}=\int_{0}^{1}\mathcal{L}(X)dX\approx\sum_{i\in\textrm{dead points}}\mathcal{L}_{i}\Delta X_{i}.$ (5)
By construction, using the likelihood value of a single dead point, $\mathcal{L}_{i}$, as a proxy for $\mathcal{L}(X)$ is an exact substitution. However, the exact values of the fractional prior volume ‘shells’, $\Delta X_{i}$, are unknown, since the exact set of shrinkage ratios, $\textbf{t}_{i}$, are unknown, and this is the dominant source of uncertainty in evidence estimation. Typically, the resulting error bar on the evidence is quantified by simulating sets of $\Delta X_{i}$ to use in the evidence calculation and quoting the standard deviation from this set of estimates.
@@ 2.3 Parameter estimation
$E[f(\theta)]=\int f(\theta)\frac{\mathcal{L}(\theta)\pi(\theta)}{\mathcal{Z}}d\theta=\frac{1}{\mathcal{Z}}\int\tilde{f}(X)\mathcal{L}(X)dX.$ (6)
$\tilde{f}(X)$ represents the prior expectation of $f(\theta)$ given that $\mathcal{L}(\theta)=\mathcal{L}(X)$:
$\tilde{f}(X)=E_{\pi|\mathcal{L}(\theta)=\mathcal{L}(X)}[f(\theta)].$ (7)
@@
$\frac{1}{\mathcal{Z}}\int\tilde{f}(X)\mathcal{L}(X)dX\approx\frac{1}{\mathcal{Z}}\sum_{i}\tilde{f}(X_{i})\mathcal{L}_{i}\Delta X_{i}.$ (8)
@@
Again, the exact prior volumes are unknown, but there is an additional source of uncertainty here which was not present before: we are required to use a single $f(\theta)$ value on each contour, $f(\theta_{i})$, as a proxy for $\tilde{f}(X_{i})$. Unlike with the likelihood, it is no longer true that this is an exact substitution, and in parameter estimation this becomes the dominant source of uncertainty. The stochasticity of nested sampling means that over a single run we will not necessarily be able to get representative values of $f(\theta)$ along every contour, and thus won’t be able to capture our true error bar over multiple runs, even when simulating sets of prior volumes. This is demonstrated in Figure 1, where we show the example of trying to estimate the sum, $f(\theta)=\theta_{1}+\theta_{2}$, of parameters in a simple two-parameter case; here, the values of $f(\theta)$ at each of the dead points on the contours do not reflect the average value over the contours. The key to capturing this uncertainty is better understanding the variation of $f(\theta)$ along the contours.
3 Methods
@@
In order to study the nested sampling parameter estimation uncertainty, we consider the example of a simulated black hole binary (BBH) signal with parameters similar to those of GW150914 and a signal-to-noise ratio (SNR) of 51. The exact injected parameter values are shown in Figure 2. Since studying the error bars carefully requires hundreds of nested sampling runs to be performed, we work within a simplified framework in which only four of the parameters are sampled: the chirp mass ($\mathcal{M}$), mass ratio ($q$), luminosity distance ($d_{L}$), and zenith angle between the total angular momentum and line of sight ($\theta_{jn}$). For the results in the following subsection, and for those in section 4 unless otherwise specified, we perform nested sampling runs with $500$ live points per run and the default PolyChord chain length of $5*\textrm{ndims}=20$. All other sampler settings were set to the PolyChord defaults. The runs are performed using bilby (Ashton et al., 2019), with a modified version of the in-built PolyChord sampler option. All waveforms are both injected and sampled with the IMRPhenomPv2 waveform model (Hannam et al., 2014) and we use two interferometers, LIGO-Hanford (H1) and LIGO-Livingston (L1).
@@
Figure 2: Posteriors recovered from the injected signal using PolyChord, with the injected parameter values indicated. Since parameter estimation studies require large numbers of runs, we only sample these 44 parameters, with the other parameters simply set to their injected values.
3.1 Evidence uncertainty estimation method is insufficient for parameter estimation
@@
Figure 3: Typically, a distribution of evidence estimates is computed from a single run using the ‘simulated weights’ method (top left). The mean and standard deviation of this distribution are then quoted as the evidence value and its corresponding uncertainty (black). If the error bars from single runs accurately quantify the uncertainty on the evidence due to unknown prior volumes, the overall mean evidence, computed over many runs, should lie in the $1^{\textrm{st}}$ percentile of estimates $1\%$ of the time, in the $50^{\textrm{th}}$ percentile $50\%$ of the time, and so on. Hence, plotting a histogram of the percentiles from each run in which the overall mean (blue dashed line) lies should give a uniform distribution from 0 to 100 (bottom left). We see here that this is indeed the case for the evidences, with the Kolmogorov-Smirnov $p$-value being 0.25, demonstrating that resampling the prior volumes is sufficient for estimating the error bar on the evidence for this example. For the chirp mass, even by eye the estimates per run are not wide enough to capture the variation in estimates between runs (top and middle right). This indicates that the variation in chirp mass along a given contour is not being properly accounted for. The percentiles plot (bottom right) confirms that the uncertainty estimate on a single nested sampling run is too optimistic. More often than would be the case for correctly distributed errors, the overall mean estimate for the chirp mass over 100 runs (blue dashed line) lies deep in the tails of the estimated chirp mass from a single run. The simulated weights method is not sufficient for calculating the parameter estimation uncertainty.
@@
The stochasticity of the fractional prior volumes in nested sampling means that the estimated mean evidence will vary from run to run. In order to accurately quantify the uncertainty on a single run, therefore, the error bar must reflect this run-to-run variation. For evidence estimation, this is typically achieved by sampling many sets of the shrinkage ratios, t, from a known distribution and using these to calculate the evidences, quoting the error bar from the spread of these estimates. Higson et al. (2018a) term this the ‘simulated weights method’ and we shall refer to it henceforth as such too.
@@
In the context of our simulated BBH signal, we can carefully verify that this method does indeed quantify the evidence uncertainty correctly. To do this, we first perform $100$ nested sampling runs on our simulated signal. For each run, we apply the simulated weights method using anesthetic (Handley, 2019) to calculate a set of evidence estimates. The resulting distributions of evidences are plotted in Figure 3. The dashed line represents the overall mean evidence calculated from all $100$ runs.
@@
The errors should be normally distributed, meaning that about $68\%$ of the time our ‘true evidence’ (estimated from the mean of $100$ runs) should lie within the error bars, and $95\%$ of the time it should lie within two times the extent of the error bars. This condition can also be viewed in terms of the distribution of the percentiles of the ‘true evidence’. If the errors are indeed correctly distributed, a histogram of these percentiles will yield a uniform distribution from 0 to $100$. For our simulated BBH example, this histogram is plotted in Figure 3, along with the Kolmogorov-Smirnov (K-S) test $p$-value for a uniform distribution. We can conclude from this that the simulated weights method leads to correct evidence uncertainty estimates from a single run.
@@
Skilling (2006) suggests using the same method to estimate parameter estimation uncertainties. As an example, we shall attempt to estimate the chirp mass of our signal. As before, for each of the $100$ runs, we can simulate sets of shrinkage ratios to obtains sets of $\Delta X$ and substitute these into equation 8 to calculate a set of chirp mass estimates. These are plotted in Figure 3, along with the overall mean chirp mass estimated from all $100$ runs.
@@
This time, even by eye it seems that the resulting distributions of chirp masses per run are not quite wide enough to capture the run-to-run variation. Performing a similar analysis of the percentiles as for the evidences yields the histogram in Figure 3. It is unmistakable from both the plot and the K-S test $p$-value that the simulated weights method alone does not lead to correct uncertainty estimates on the chirp mass. Far more often than ought to be the case, the mean chirp mass over many runs lies deep in the tails of the estimates from a single run. This is an empirical verification, in the context of gravitational waves, of results already established theoretically by Chopin & Robert (2010) and experimentally by Higson et al. (2018a).
@@ 3.2 Phantom points
The key piece lacking from the simulated weights method is that it uses the single value of $f(\theta)$ available for each iso-likelihood contour as a proxy for the average $f(\theta)$ over the contour, since there is no extra information in the live points to do otherwise. However, phantom points, drawn from the likelihood-constrained prior and each with a corresponding likelihood evaluation, have the potential to provide this additional information about the variation of $f(\theta)$ along contours. Below, we present two novel uncertainty estimation methods which incorporate this.
3.3 Phantom points-informed uncertainty estimation
3.3.1 Method 1: Likelihood binning